find the slope of (0,0) (-6,2)
Answer by sdmmadam@yahoo.com(530)

find the slope of line joining (0,0)and (-6,2)
[The slope of the line joining two given points
A(x1,y1) and B(x2,y2) is given by
m = (y2-y1)/(x2-x1)
Here in the given problem
A(x1,y1) is the origin O(0,0) Therefore x1 = 0, y1 = 0
and B(x2,y2) is the (-6,2) Therefore x2 = -6, y2 = 2]
the slope of line joining (0,0)and (-6,2)is
m = (2-0)/[(-6)-0]
=2/(-6)
=(-1/3)